OpenStudy (anonymous):

best possible first step to solving the system by first eliminating the x variable. 3x + 5y = 1 2x + 4y = -4 do not know what I am doing

OpenStudy (anonymous):

choices are Multiply the first equation by 2 and multiply the second equation by 3. Multiply the first equation by -2 and multiply the second equation by 3. Multiply the first equation by -2 and multiply the second equation by -3. None of the above

OpenStudy (shinalcantara):

the first three choices can be correct depending on what process you'll involve the equations with to eliminate. say: (3x +5y=1)*2 (2x + 4y =-4) *3 you can eliminate x by subtracting the 2 equations or (3x +5y=1)*-2 (2x + 4y =-4) *3 you can eliminate by adding both equations or (3x +5y=1)*-2 (2x + 4y =-4) *-3 you can eliminate by subtracting the two equations

OpenStudy (anonymous):

thank you